# Volume and topology of bounded and closed hyperbolic $3$-manifolds, II

### v1

@@ 6.2. Background on least-area surfaces
[…]
@@
If (${}^{\prime}$) holds, i.e. if $f$ is an embedding, then it follows from 6.2.7.4 that Alternative (i-b) of the present proposition holds.
Now suppose that (${}^{\prime}$) holds. Thus $f$ is a two-sheeted covering map to a one-sided surface $K\subset DN$. According to 6.2.7.4 we have $K\subset\mathop{\rm int}N$. Let $J$ be a tubular neighborhood of $K$ in $\mathop{\rm int}N$. Then $J$ is a twisted $I$-bundle with core $K$. Now since $S$ is orientable and $K$ is not, the covering map $f:S\to K$ must be equivalent to the orientation covering of $K$. It follows that $f$ is homotopic in $J\subset\mathop{\rm int}N$ to a diffeomorphism $g:S\to\partial J$. But by 6.2.7.4, $f$ is homotopic in $\mathrm{int}\,N$ to the inclusion $S\hookrightarrow N$. Hence the maps $g$ and $S\hookrightarrow N$, which may both be regarded as embeddings of $S$ in $\mathop{\rm int}N$, are homotopic in $\mathop{\rm int}N$. It now follows from Corollary 5.5 of [36] that the surfaces $S$ and $\partial J=g(S)$ are isotopic in $N$.
@@
Let $S$ be a connected incompressible surface in a compact, orientable hyperbolic $3$-manifold $M$ with empty or connected, totally geodesic boundary (so that each component of $M\,\backslash\backslash\,S$ has non-empty boundary and is simple by the discussion in 6.1, and hence $\mathop{\rm kish}(M\,\backslash\backslash\,S)$ is defined). Then we have
@@ 7. The capstone volume-topology dichotomy in the geodesic boundary case
As recorded in the proof of [14, Th. 7.4], the condition 7.1.1 implies that $T=\partial V$ is incompressible in $N$. According to the discussion in Subsection 6.1 of the present paper, this implies that the components of $N\,\backslash\backslash\,T$, which are canonically identified with $X$ and $V$, are simple; they obviously have non-empty boundaries, and so $\mathop{\rm kish}V$ and $\mathop{\rm kish}X$ are defined. It is also recorded in the proof of [14, Th. 7.4] that $\mathop{\rm kish}V$ and $\mathop{\rm kish}X$ are non-empty. Therefore $\bar{\chi}(\mathop{\rm kish}(N\,\backslash\backslash\,T))\geq 2$, and applying Theorem 6.2.8 with $T$ playing the role of $S$ there gives the desired volume bound. ∎
@@ 8. Background for the results on closed manifolds
Let $k\geq 3$ be an integer, and let $M$ be a closed, orientable, hyperbolic $3$-manifold. Suppose that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$, and that $\pi_{1}(M)$ is not $k$-free. Then $M$ contains a closed incompressible surface of some genus $g$ with $12\leq g\leq k-$ 1.
@@
Proposition 8.1 of [11] includes the fact that if $k\geq 3$ is an integer and $M$ is a closed, orientable hyperbolic $3$-manifold with $H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$, then either $\pi_{1}(M)$ is $k$-free, or M contains a closed incompressible surface of genus at most $k-1$. Since the genus of an incompressible surface in $M$ is at least $2$ by 8.1, the present proposition follows. ∎
@@ 9. Incompressible surfaces, homology rank, and volume
This is a formal consequence of Propositions 6.1 and 6.2 of [8]. Note that if $M$, $S$ and $A$ satisfy the hypotheses of Proposition 9.3, then $A$ is irreducible, boundary-irreducible and acylindrical. In Section 6 of [8], a certain real-valued invariant $\mathop{\rm geodvol}A$ is defined for a compact, connected $3$-manifold $A$ with non-empty boundary. Proposition 6.1 of [8] asserts that if $A$ is irreducible, boundary-irreducible and acylindrical, then $A$ is diffeomorphic to a hyperbolic manifold $N$ with totally geodesic boundary, and $\mathop{\rm vol}N=\mathop{\rm geodvol}A$. Proposition 6.2 of [8] asserts that if $A$ is an acylindrical component of $M\setminus\setminus S$, where $S$ is a connected, incompressible surface in a closed, orientable hyperbolic $3$-manifold $M$, then $\mathop{\rm vol}M\geq\mathop{\rm geodvol}A$. Proposition 9.3 now follows immediately.
@@
We shall assume that $\pi_{1}(M)$ is not $k$-free, and show that $\mathop{\rm vol}M>2V_{\rm oct}$. Since $\pi_{1}(M)$ is not $k$-free and $r\geq\max(3k-4,6)$, it follows from Proposition 8.3 that $M$ contains a closed incompressible surface of some genus $g$ with $12\leq g\leq k-$ 1.
@@
Fix an acylindrical component $A$ of $M\setminus\setminus S$, and let $B$ denote the other component of $M\setminus\setminus S$. Since $A$ is acylindrical, it follows from Proposition 9.3 that $A$ is diffeomorphic to a hyperbolic $3$-manifold $N$ with totally geodesic boundary, and that
@@
[…]
@@ 10. Homology of manifolds with volume at most $V_{\rm oct}/2$
If $\Pi$ is $3$-free, it follows from Proposition 8.4 that $\mathop{\rm vol}M>3.08$. Since $3.08>V_{\rm oct}/2\geq\mu$, the conclusion holds in this case. For the rest of the proof we shall assume that $\Pi$ is not $3$-free. We fix a subgroup $\Delta$ of $\Pi$ which has rank at most $3$ and is not free. The image $J$ of the inclusion homomorphism $H_{1}(\Delta;{\mathbb{F}}_{2})\to H_{1}(\Pi;{\mathbb{F}}_{2})=V$ has dimension at most $3$ since $\mathop{{\rm rank}}\Delta\leq 3$. Since $\dim V\geq 5$, there is a codimension-$2$ subspace $P$ of $V$ containing $J$. The subspace $P$ defines a regular covering space $\widetilde{M}$ of $M$ whose covering group is isomorphic to ${\mathbb{Z}}/2{\mathbb{Z}}\times{\mathbb{Z}}/2{\mathbb{Z}}$. Since $P\supset J$, there is a subgroup of $\pi_{1}(\widetilde{M})$ isomorphic to $\Delta$. Hence $\pi_{1}(\widetilde{M})$ is not $3$-free.
We have $\dim P=\dim V-2\geq 3$. Hence if we set ${\widetilde{r}}=\dim H_{1}(\widetilde{M};{\mathbb{F}}_{p})$, then the case $p=2$ of Lemma 8.5 gives ${\widetilde{r}}\geq 2\cdot 3+1=7$. Since in particular we have ${\widetilde{r}}\geq 6$, and $\pi_{1}(\widetilde{M})$ is not $3$-free, we may apply Proposition 9.4, with $k=3$ and with $\widetilde{M}$ playing the role of $M$, to deduce that $\mathop{\rm vol}\widetilde{M}>2V_{\rm oct}$. Since $\widetilde{M}$ is a four-fold covering of $M$, we have $\mathop{\rm vol}M=(\mathop{\rm vol}\widetilde{M})/4>V_{\rm oct}/2$, as required. ∎
@@ 11. Volumes of manifolds with small cup product rank
Suppose that $M$ is a closed, aspherical $3$–manifold. Set $r=\dim H_{1}(M;{\mathbb{F}}_{2})$, and let $t$ denote the dimension of the image of the cup product pairing $H^{1}(M;{\mathbb{F}}_{2})\otimes H^{1}(M;{\mathbb{F}}_{2})\to H^{2}(M;{\mathbb{F}}_{2})$. Then for any two-sheeted covering $\widetilde{M}$ of $M$, we have $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 2r-t-1$.
@@
We now prove both assertions in the case where $\pi_{1}(M)$ is not $4$-free. In this case we fix a subgroup $\Delta$ of $\Pi$ which has rank at most $4$ and is not free. The image $J$ of the inclusion homomorphism $H_{1}(\Delta;{\mathbb{F}}_{2})\to H_{1}(\Pi;{\mathbb{F}}_{2})=V$ has dimension at most $4$ since $\mathop{{\rm rank}}\Delta\leq 4$. Since the hypothesis of either of the assertions (1) or (2) implies $\dim V\geq 5$, there is a codimension-$1$ subspace $P$ of $V$ containing $J$. The subspace $P$ defines a two-sheeted covering space $\widetilde{M}$ of $M$. Since $P\supset J$, there is a subgroup of $\pi_{1}(\widetilde{M})$ isomorphic to $\Delta$. Hence $\pi_{1}(\widetilde{M})$ is not $4$-free.
According to Lemma 11.1, we have $\dim H_{1}(\widetilde{M};{\mathbb{F}}_{2})\geq 2r-t-1$. But the hypothesis of either of the assertions (1) or (2) implies that $2r-t-1\geq 8$, and hence $\dim H_{1}(\widetilde{M};{\mathbb{F}}_{2})\geq 8$. Since $\pi_{1}(\widetilde{M})$ is not $4$-free, we may now apply Proposition 9.4, taking $k=4$, and letting $\widetilde{M}$ play the role of $M$, to deduce that $\mathop{\rm vol}\widetilde{M}>2V_{\rm oct}$. Since $\widetilde{M}$ is a two-sheeted covering of $M$, we have $\mathop{\rm vol}\widetilde{M}=(\mathop{\rm vol}M)/2>V_{\rm oct}$. Since $V_{\rm oct}>3.57$, this establishes both Assertion (1) and Assertion (2) in this case. ∎

### v2

@@ 6.2. Background on least-area surfaces
To prove 6.2.7.4, we first observe that the inclusion map $S\hookrightarrow N$ lifts to an embedding $j$ of $S$ in $\widetilde{DN}$. By the covering homotopy property of covering spaces, $j$ is homotopic in $\widetilde{DN}$ to some lift $\widetilde{f}$ of $f$. It follows from (6.2.7.2) that either $\widetilde{f}(S)\subset\mathop{\rm int}N_{0}$ or $\widetilde{f}(S)\subset\mathop{\rm int}Z$. But if $\widetilde{f}(S)$ were contained in $\mathop{\rm int}Z$, then by 6.2.7.3, $\widetilde{f}$ would be homotopic in $\widetilde{DN}$ to a map with image contained in $\partial N_{0}$, and hence $S\hookrightarrow\mathit{DN}$ would be homotopic in $\mathit{DN}$ to a map with image contained in $\partial N$; this contradicts 6.2.7.1. It follows that $\widetilde{f}(S)\subset\mathop{\rm int}N_{0}$. Now since $j$ and $\widetilde{f}$ are homotopic in $\widetilde{DN}$ and both map $S$ into $\mathop{\rm int}N_{0}$, and since we have observed that $N_{0}\hookrightarrow\widetilde{DN}$ is a homotopy equivalence, the maps $j$ and $\widetilde{f}$ are in fact homotopic in $\mathop{\rm int}N_{0}$. This immediately implies 6.2.7.4.
@@
If (i) holds, i.e. if $f$ is an embedding, then it follows from 6.2.7.4 that Alternative (i-b) of the present proposition holds.
Now suppose that (ii) holds. Thus $f$ is a two-sheeted covering map to a one-sided surface $K\subset DN$. According to 6.2.7.4 we have $K\subset\mathop{\rm int}N$. Let $J$ be a tubular neighborhood of $K$ in $\mathop{\rm int}N$. Then $J$ is a twisted $I$-bundle with core $K$. Now since $S$ is orientable and $K$ is not, the covering map $f:S\to K$ must be equivalent to the orientation covering of $K$. It follows that $f$ is homotopic in $J\subset\mathop{\rm int}N$ to a diffeomorphism $g:S\to\partial J$. But by 6.2.7.4, $f$ is homotopic in $\mathrm{int}\,N$ to the inclusion $S\hookrightarrow N$. Hence the maps $g$ and $S\hookrightarrow N$, which may both be regarded as embeddings of $S$ in $\mathop{\rm int}N$, are homotopic in $\mathop{\rm int}N$. It now follows from Corollary 5.5 of [36] that the surfaces $S$ and $\partial J=g(S)$ are isotopic in $N$.
@@
[…]
@@ 7. The capstone volume-topology dichotomy in the geodesic boundary case
[…]
@@ 8. Background for the results on closed manifolds
Let $k\geq 3$ be an integer, and let $M$ be a closed, orientable, hyperbolic $3$-manifold. Suppose that $\dim H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$, and that $\pi_{1}(M)$ is not $k$-free. Then $M$ contains a closed incompressible surface of some genus $g$ with $2\leq g\leq k-1$.
@@
Proposition 8.1 of [11] includes the fact that if $k\geq 3$ is an integer and $M$ is a closed, orientable hyperbolic $3$-manifold with $H_{1}(M;{\mathbb{F}}_{2})\geq\max(3k-4,6)$, then either $\pi_{1}(M)$ is $k$-free, or M contains a closed incompressible surface of genus at most k1k-1. Since the genus of an incompressible surface in $M$ is at least $2$ by 8.1, the present proposition follows. ∎
@@ 9. Incompressible surfaces, homology rank, and volume
This is a formal consequence of Propositions 6.1 and 6.2 of [8]. Note that if $M$, $S$ and $A$ satisfy the hypotheses of Proposition 9.3, then $A$ is irreducible, boundary-irreducible and acylindrical. In Section 6 of [8], a certain real-valued invariant $\mathop{\rm geodvol}A$ is defined for a compact, connected $3$-manifold AA with non-empty boundary. Proposition 6.1 of [8] asserts that if $A$ is irreducible, boundary-irreducible and acylindrical, then $A$ is diffeomorphic to a hyperbolic manifold $N$ with totally geodesic boundary, and $\mathop{\rm vol}N=\mathop{\rm geodvol}A$. Proposition 6.2 of [8] asserts that if $A$ is an acylindrical component of $M\setminus\setminus S$, where $S$ is a connected, incompressible surface in a closed, orientable hyperbolic $3$-manifold $M$, then $\mathop{\rm vol}M\geq\mathop{\rm geodvol}A$. Proposition 9.3 now follows immediately.
@@
We shall assume that $\pi_{1}(M)$ is not $k$-free, and show that $\mathop{\rm vol}M>2V_{\rm oct}$. Since $\pi_{1}(M)$ is not $k$-free and $r\geq\max(3k-4,6)$, it follows from Proposition 8.3 that $M$ contains a closed incompressible surface of some genus $g$ with $2\leq g\leq k-1$.
@@
Fix an acylindrical component $A$ of $M\setminus\setminus S$, and let $B$ denote the other component of $M\setminus\setminus S$. Since $A$ is acylindrical, it follows from Proposition 9.3 that $A$ is diffeomorphic to a hyperbolic $3$-manifold NN with totally geodesic boundary, and that
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${J\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j}$${H_{1}(A)\oplus H_{1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H_{1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
@@ 10. Homology of manifolds with volume at most $V_{\rm oct}/2$
[…]
@@ 11. Volumes of manifolds with small cup product rank
Suppose that $M$ is a closed, aspherical 33–manifold. Set $r=\dim H_{1}(M;{\mathbb{F}}_{2})$, and let $t$ denote the dimension of the image of the cup product pairing $H^{1}(M;{\mathbb{F}}_{2})\otimes H^{1}(M;{\mathbb{F}}_{2})\to H^{2}(M;{\mathbb{F}}_{2})$. Then for any two-sheeted covering $\widetilde{M}$ of $M$, we have $\dim H_{1}(M;{\mathbb{F}}_{2})\geq 2r-t-1$.
@@
[…]